find the interest: deposit $800, rate is %6 per year , time is 2 years
Found 2 solutions by Greenfinch, checkley77:
Answer by Greenfinch(383)

800 x 1.06^2 = $ 898.88

Answer by checkley77(12844)

800(1+.06)^2
800(1.06)^2
800*1.1236=898.88
898.88-800=$98.88 is the interest for the 2 year deposit @ 6%.